OpenStudy (anonymous):

mary races her bicycle for 128m. a wheel of her bicycle 32 times as the bycicle travels that distance. what is the diameter of the wheel?

OpenStudy (anonymous):

The perimeter of a circle is: \[2 \pi r\] You're "laying down" 32 perimeters to equal 128 meters.

OpenStudy (anonymous):

so 2 time 3.14 times 32?

OpenStudy (anonymous):

Yes, that will equal 128 meters.

OpenStudy (anonymous):

i got 201

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

\[32(2 \pi r) = 128\] \[2 \pi r = 4\]